OpenStudy (anonymous):

how many molecules of C6H1206 contain 4.56 x 10^23 atoms of carbon?

OpenStudy (anonymous):

\[4.56x10^23 C atom (\frac{1 mole C atom}{6.022x10^23 C atom})(\frac{1 mole C_6H_12O_6}{6 mole C atom})(\frac{6.022x10^23 C_6H_12O_6 molec}{1 mole C_6H_12O_6})=\]

OpenStudy (anonymous):

so basically its 4.56x10^32 divided by 6 and i get 7.60x10^31 molecules as the answer?

OpenStudy (anonymous):

I think that number is to big

OpenStudy (anonymous):

7.6x10^23 is the answer i got